OpenStudy (aravindg):

If f(x)= x2 + 3 x + 2, find that value of x for which f(x + 1) = f(x).

OpenStudy (anonymous):

(x+1)^2 +3(x+1)+2=x^2+3x+2

OpenStudy (anonymous):

(x+1)^2 +3(x+1)+2=x^2+3x+2 x=-2

OpenStudy (anonymous):

2x+4=0

OpenStudy (aravindg):

wow thx